OpenStudy (anonymous):

Solve. 11^x – 1 = 11^5x –19

OpenStudy (ash2326):

We have \[11^{x-1}=11^{5x-19}\] equate the powers both sides \[x-1=5x-19\] swapping the sides \[5x-19=x-1\] Add 19 to both sides \[5x=x-1+19\] subtract both sides by x \[5x-x=18\] or \[4x=18\] we get \[x=4.5 or \frac{9}{2}\]